# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

### v2

@@ Abstract
Considering the challenges faced by large language models (LLMs) in logical reasoning and planning, prior efforts have sought to augment LLMs with access to external solvers. While progress has been made on simple reasoning problems, solving classical constraint satisfaction problems, such as the Boolean Satisfiability Problem (SAT) and Graph Coloring Problem (GCP), remains difficult for off-the-shelf solvers due to their intricate expressions and exponential search spaces. In this paper, we propose a novel differential logic layer-aided language modeling (DiLA) approach, where logical constraints are integrated into the forward and backward passes of a network layer, to provide another option for LLM tool learning. In DiLA, LLM aims to transform the language description to logic constraints and identify initial solutions of the highest quality, while the differential logic layer focuses on iteratively refining the LLM-prompted solution. Leveraging the logic layer as a bridge, DiLA enhances the logical reasoning ability of LLMs on a range of reasoning problems encoded by Boolean variables, guaranteeing the efficiency and correctness of the solution process. We evaluate the performance of DiLA on two classic reasoning problems and empirically demonstrate its consistent outperformance against existing prompt-based and solver-aided approaches.
@@ 1 Introduction
Recently, a significant research thrust has been on leveraging large language models (LLMs) for reasoning and planning, with numerous efforts aimed at augmenting their reasoning capabilities. These endeavors include text-based temporal reasoning Xiong et al. (2024), logic feedback-enhanced alignment methods Nguyen et al. (2023), and prompt-based methods such as the chain of thoughts Wang et al. (2023a) or a simple directive like “Let’s think step by step” Lightman et al. (2023). Despite the strides made by LLMs in achieving human-like reasoning abilities, they still encounter challenges when confronted with complex plannings Valmeekam et al. (2022). Occasionally, LLMs exhibit unfaithful reasoning, leading to derived conclusions that do not consistently follow the previously generated reasoning chain in practical applications Pan et al. (2023).
To this end, recent works have begun to augment LLMs with access to external solvers, i.e., utilize LLMs to first parse natural language logical questions into symbolic representations and subsequently employ external solvers to generate answers based on these representations. To enhance parsing accuracy, LoGiPT Feng et al. (2023) has been proposed to directly emulate the reasoning processes and mitigate the parsing errors by learning to strict adherence to solver syntax and grammar. It is fine-tuned on a constructed instruction-tuning dataset derived from revealing and refining the invisible reasoning process of deductive solvers. An alternative approach for improving reasoning capabilities involves SATLM, a new satisfiability-aided modeling approach Ye X (2023), in which an LLM is used to generate a declarative task specification with few-shot prompting and offload the actual reasoning task to an off-the-shelf theorem prover, i.e., Z3 solver De Moura and Bjørner (2008). This method has demonstrated superiority over program-aided Language Models Gao et al. (2023) by achieving a $23\%$ improvement on the GSM arithmetic reasoning dataset, establishing a new state-of-the-art.
@@
Despite their impressive performance on various benchmark tests, SATLM Ye X (2023) and its solver-augmented successors can not deal with reasoning and planning problems in practical scenarios. Particularly, these real problems often state a set of premises and complex constraints and require a sophisticated search process to find the optimal solution, which is still challenging even for modern solvers. More specifically, even the state-of-the-art symbolic solvers, such as Z3De Moura and Bjørner (2008) and KissatBiere and Fleury (2022), encounter significant bottlenecks when addressing formal verification problems such as electronic circuit verification involving tree/cyclic circuit structuresShi et al. (2023) and job-shop scheduling problems Li et al. (2022) with complex graph connections. For instance, the Kissat solver requires days or even weeks to process a vanilla circuit verification problem with arithmetic circuit modules including multipliers or multiply-add circuits, yet fails to produce a solution. This highlights the limitations of solver-augmented LLM approaches which solely rely on symbolic solvers to tackle logical reasoning problems in reality.
@@
  • Leveraging SAT encoding as a bridge, DiLA successfully translates natural language reasoning problems into satisfiability problems, enabling it to tackle a range of constraint satisfaction problems, such as Boolean satisfiability and graph coloring.
@@
We evaluate the performance of our approach on two classical reasoning problems: Boolean satisfiability and graph coloring. Our analysis yields two key findings: firstly, on simple artificial problem instances, DiLA boosts the inference accuracy of LLMs to 100% and consistently outperforms solver-aided approaches with improved runtime. Secondly, for real-world problems that current solvers struggle with, DiLA showcases robustness and remarkable efficiency in handling these complex test cases, thereby opening up opportunities for further real-world applications.
@@ 2 Motivation
This paper explores the potential of language understanding and logical reasoning capabilities in LLMs. Traditionally, the prevailing approaches have either relied solely on LLMs for step-by-step reasoning Wang et al. (2023a); Zhou et al. (2022) or offloaded reasoning tasks to off-the-shelf solvers Feng et al. (2023); Ye X (2023). However, we propose a third approach, as these two extremes either underutilize or over-rely on LLMs during reasoning. Specifically, DiLA leverages the powerful understanding abilities of LLMs to extract logical formulas and generate possible solutions. By utilizing LLMs, such as GPT-4, as a solution generation engine, we can produce an initial solution based on input semantic constraints, laying the groundwork for further refinement.
@@
Figure 2: Exemplar comparison of solving graph coloring problems by different approaches. Direct prompt by GPT-4 makes errors when generating the color assignment step by step; SATLM based on the Z3 solver cannot solve parsed constraints and outputs unknown (both errors are highlighted in red). In contrast, DiLA can generate the correct answer by combining the strengths of LLMs and the differential logic layer.
Consequently, our objective is to identify a synergistic approach that combines the strengths of LLMs and a differential logic layer, thereby circumventing the limitations of traditional solvers. This approach leverages the LLM’s capacity to comprehend logical formulas while concurrently utilizing the logic layer’s refinement abilities to achieve accurate solutions. Figure 2 illustrates a comparison between COT, SATLM, and our proposed DiLA. The LLM alone may introduce logical flaws during step-by-step inference, such as assigning the same color to vertex 1 and vertex 4 despite their edge connection, and SATLM may struggle with complex reasoning problems due to its backbone solver’s limitations. In contrast, DiLA produces accurate answers through the collaboration of the LLM and the differential logic layer. Specifically, DiLA first uses the LLM to parse a natural language input into logic constraints and generate an initial solution based on its semantic understanding, then employs the differential logic layer to refine this initial solution.
3 Overview
In this section, we present DiLA, which augments LLM with the ability of logical reasoning by incorporating a differential logic layer. More specifically, DiLA addresses the challenge of using LLMs to tackle canonical reasoning tasks expressed in natural language, such as the Graph Coloring problem and Boolean Satisfiability problem. These tasks typically involve presenting a set of premises and constraints, prompting questions that necessitate intricate deductive reasoning over the provided inputs, which remains a formidable challenge even for contemporary LLMs Valmeekam et al. (2022). Notably, the Satisfiability (SAT) problem can serve as a versatile intermediate step for solving a broad range of constraint satisfaction problems, provided they can be expressed using Boolean variables. Specifically, problem instances from NP-complete domains, such as Graph Coloring and Set Cover, can be seamlessly encoded into SAT problem specifications, thereby allowing for efficient solutions via SAT algorithms Stechly et al. (2024); Liu et al. (2023). Therefore, we harness SAT encoding as a general bridge to tackle these constraint satisfaction problems in practical settings.
At a high level, given the natural language description of a collection of facts $\phi$ (such as propositions or constraints) about some variables $V$, the LLM is prompted with a question $Q$ related to these variables. The ultimate goal of the reasoning task is to determine an answer to $Q$, which essentially entails identifying the assignment for the variable of interest that satisfies all the input constraints $\phi$.
@@ 3 Overview
4 Differenial Logic Layer-Aided Language Models
@@ 4.1 Problem Formulator and Initialization
Intuitively, LLMs may struggle with directly solving complex reasoning problems. However, they have demonstrated a notable ability to comprehend textual inputs and translate them into formal programs, such as mathematical equations He-Yueya et al. (2023) or satisfiability modulos Ye X (2023). Therefore, we leverage this transformation capability of LLMs to formulate problems into Boolean formulas. Specifically, given a problem description in natural language, DiLA prompts an LLM with detailed instructions to generate the SAT specification, which includes a set of premises and constraints. Typically, the SAT specification here involves conjunctive normal forms (CNFs), denoted as $\phi(v_{1},\dots,v_{n})$, which is a conjunction of clauses (constraints) $C$. Formally, rules are written in the conjunctive form of clauses $C_{1}\wedge C_{2}\cdots\wedge C_{m}$, which each $C_{j}$ is a constraint. The total rule is satisfied if and only if all of the clauses $C_{1},C_{2},\dots,C_{m}$ are simultaneously True. Each clause represents a disjunction of literals, where a literal is either a propositional variable $v_{i}$ or its complement $\neg v_{i}$, for example, $v_{1}\vee\neg v_{2}$. In other words, if at least one literal in a clause is True, this clause would also be True. Variables can be assigned logic values, either $1$ or $-1$, representing True or False, respectively  [^1] In other works, they may claim the logic value of each literal is 00 or $1$. It should be noted that the two claims are equal under simple mathematical transformations.. See Appendix A for more details on the SAT specification.
Aside from problem formulation, leveraging an LLM can be a valuable strategy to generate an initial solution. We observe that, after generating the SAT specification, we can prompt LLMs, like GPT-4 Achiam et al. (2023) or Llama-3 Meta (2024), to produce a potential solution or a set of possible solutions. This can be achieved by framing the problem as a natural language query, such as “What is the logical solution based on the premises?” or “Can you provide a possible answer from these given constraints?” (see Appendix B for more details). The LLM’s response can then serve as a starting point for further refinement and validation, allowing us to build upon its output and iteratively improve the solution through logical analysis and reasoning. By harnessing the LLM’s ability to understand semantic constraints, we can tap into its potential to facilitate the initial solution-finding process and accelerate solving progress towards a well-reasoned answer.
@@ 4.2 From SAT to Differential MaxSAT Formulations
In fact, SAT is very efficient at expressing constraint satisfaction problems in that many of the standard NP-complete problems, like graph coloring, can be straightforwardly formulated and efficiently solved as SAT problems Buss and Nordström (2021); Liu et al. (2023). Traditionally, a SAT solver, such as Z3 and Kissat, is leveraged to determine a satisfying assignment for the given constraint formula $\phi$. Contemporary SAT solvers are founded on the Conflict-Driven-Clause-Learning (CDCL) algorithm, which excels in its ability to learn from conflicts and use that conflict knowledge to prune the branch-and-bound search space more effectively. However, existing CDCL-based SAT solvers still suffer from exponential searching space and are unable to correct errors through a learning-from- mistakes system, resulting in an infinite loop in solving complex SAT problems Shi et al. (2023).
In this study, when an NL reasoning problem is fed into DiLA, our goal is to determine its solution that can satisfy all logical constraints. To achieve this, we might incorporate a solver as an additional logic reasoner for the LLM, as in SATLM Ye X (2023) or LOGIC-LM Pan et al. (2023). However, the exponential search complexity inherent in heuristic-based SAT solvers poses a significant challenge, limiting their effectiveness when dealing with complex real-world problems. Therefore, a key issue is how to design an efficient solver surrogate that can both be seamlessly integrated into LLMs and efficiently address logical reasoning problems.
@@
where $\bigwedge$ is the logical “and” symbol. Indeed, the objective value in Equation 2 is 0 if and only if a satisfiable solution can be found. Our goal is to establish a continuous upper bound, referred to as the “loss”, for each clause to quantify its level of unsatisfiability. In essence, the loss takes an upper bound if the clause is unsatisfied, and by minimizing this loss, we can strive to push it closer to satisfaction. To make a purely quadratic loss function as in  Wang and Kolter (2019), we introduce $v_{0}=1$ and $s_{0j}=-1$ in Equation 2. Therefore, the minimization problem in Equation 2 can be solved by transforming into a quadratic loss function as:
@@
[Uncaptioned image] Figure 4: Examplar illustration of VCG. Algorithm 1 The forward pass of logic layer 0:  Solution $\boldsymbol{v}^{k}\in\mathbb{R}^{n}$ at $k$-th epoch. 0:  $y^{k}$, final solution $\boldsymbol{v}^{*}$. 1:  $\tilde{\boldsymbol{v}}^{k}\leftarrow[\boldsymbol{v}^{k}_{i}>0:i=1,\dots,n]$; 2:  $\phi^{\prime}\leftarrow\phi(\tilde{v}^{k}_{1},\dots,\tilde{v}^{k}_{n})$; 3:  if $\phi^{\prime}$ is satisfiable then 4:     $y^{k}\leftarrow$ True; 5:     $\boldsymbol{v}^{*}\leftarrow\tilde{\boldsymbol{v}}^{k}$ ; 6:  else 7:     $y^{k}\leftarrow$ False; 8:  end if
@@ 4.3 Differential Logic Layer
We envision the logic layer being used primarily at the top of LLMs, embedding logical formulas $\phi$ produced by the backbone LLM, taking LLM-initialized variable assignments as inputs, and producing outputs that are consistent with $\phi$. Specifically, we draw an analogy between the fully connected (FC) layer and the variable-clause graph (VCG), as depicted in Section 4.2, where the weight for positive literals in the clause is 1 (solid arrow) and negative literals in the clause is -1 (dashed arrow). Specifically, we map each variable to an input neuron in an FC layer, each clause to an output neuron, and the coefficients to the weights in the linear transformation with a zero bias vector. While this analogy holds, there are two crucial differences between traditional FC layers and our proposed logic layer. Firstly, the differential logic layer has no unknown parameters, whereas FC layers require data-driven training to learn their weights. Secondly, each clause in the logic layer is only partially connected to variables, in contrast to fully connected layers, where all input neurons are connected to all output neurons.
@@
Backward Pass. The backward pass is responsible for computing the gradients of the layer inputs and derives updates to variables that steer towards satisfying the constraints $\phi$. A crucial aspect of the backward pass is identifying the input variables that contribute most to the unsatisfiability of the constraint formulas. It is well-established that variables in $\bar{I}$ form the unsatisfiable subset and are, therefore, more likely to be sources of conflict. Conversely, variables not present in $\bar{I}$ can have their gradients set to zero, as their absence in the conflict clauses provides no evidence regarding the correctness or incorrectness of their values. Inspired by the stochastic local search (SLS) algorithm, commonly used in constraint satisfaction problems Chu et al. (2023), we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value at each iteration. However, our logic layer diverges from SLS in that it employs a “differential” variable selection mechanism during backpropagation, whereas SLS relies on meta-heuristics.

### v3

@@ Abstract
Considering the challenges faced by large language models (LLMs) in logical reasoning and planning, prior efforts have sought to augment LLMs with access to external solvers. While progress has been made on simple reasoning problems, solving classical constraint satisfaction problems, such as the Boolean Satisfiability Problem (SAT) and Graph Coloring Problem (GCP), remains difficult for off-the-shelf solvers due to their intricate expressions and exponential search spaces. In this paper, we propose a novel differential logic layer-aided language modeling (DiLA) approach, where logical constraints are integrated into the forward and backward passes of a network layer, providing another option for LLM tool learning. In DiLA, LLM aims to transform the language description to logic constraints and identify initial solutions of the highest quality, while the differential logic layer focuses on iteratively refining the LLM-prompted solution. Leveraging the logic layer as a bridge, DiLA enhances the logical reasoning ability of LLMs on a range of reasoning problems encoded by Boolean variables, guaranteeing the efficiency and correctness of the solution process. We evaluate the performance of DiLA on three classic constraint satisfaction problems and empirically demonstrate its consistent outperformance against existing prompt-based and solver-aided approaches.
DiLA: Enhancing LLM Tool Learning with Differential Logic Layer
Yu Zhang1  Hui-Ling Zhan2  Zehua Pei1  Yingzhao Lian2  Lihao Yin2 Mingxuan Yuan2Bei Yu1 1The Chinese University of Hong Kong  2Noah’s Ark Lab, Huawei
@@ 1 Introduction
Recently, a significant research thrust has been on leveraging large language models (LLMs) for reasoning and planning, with numerous efforts aimed at augmenting their reasoning capabilities. These endeavors include text-based temporal reasoning (Xiong et al., 2024), logic feedback-enhanced alignment methods (Nguyen et al., 2023), and prompt-based methods such as the chain of thoughts (CoT) (Wang et al., 2023) or a simple directive like “Let’s think step by step” (Lightman et al., 2023). Despite the strides made by LLMs in achieving human-like reasoning abilities, they still encounter challenges when confronted with complex plannings (Valmeekam et al., 2022). Occasionally, LLMs exhibit unfaithful reasoning, leading to derived conclusions that do not consistently follow the previously generated reasoning chain in practical applications (Pan et al., 2023).
To this end, recent works have begun to augment LLMs with access to external solvers, i.e., utilize LLMs to first parse natural language logical questions into symbolic representations and subsequently employ external solvers to generate answers based on these representations. To enhance parsing accuracy, LoGiPT (Feng et al., 2023) has been proposed to directly emulate the reasoning processes and mitigate the parsing errors by learning to strict adherence to solver syntax and grammar. It is fine-tuned on a constructed instruction-tuning dataset derived from revealing and refining the invisible reasoning process of deductive solvers. An alternative approach for improving reasoning capabilities involves LOGIC-LM, a new symbolic solver-aided modeling approach (Pan et al., 2023), in which an LLM is used to generate a symbolic problem formulation with in-context learning and offload the actual reasoning task to off-the-shelf symbolic reasoners, such as Z3 solver (De Moura and Bjørner, 2008). This solver-augmented method has demonstrated superiority over CoT by achieving a large margin of $18.4\%$ on average, establishing a new state-of-the-art.
@@
Despite their impressive performance on various benchmark tests, LOGIC-LM (Pan et al., 2023) and its solver-augmented successors can not deal with reasoning and planning problems in practical scenarios. Particularly, these real problems often state a set of premises and complex constraints and require a sophisticated search process to find the optimal solution, which is still challenging even for modern solvers. More specifically, even the state-of-the-art symbolic solvers, such as Z3(De Moura and Bjørner, 2008) and Kissat(Biere and Fleury, 2022), encounter significant bottlenecks when addressing formal verification problems such as electronic circuit verification involving tree/cyclic circuit structures(Shi et al., 2023) and job-shop scheduling problems (Li et al., 2022) with complex graph connections. For instance, the Kissat solver requires days or even weeks to process a vanilla circuit verification problem with arithmetic circuit modules including multipliers or multiply-add circuits, yet fails to produce a solution. This highlights the limitations of solver-augmented LLM approaches which solely rely on symbolic solvers to tackle logical reasoning problems in reality.
@@
  • Leveraging SAT encoding as a bridge, DiLA successfully translates natural language reasoning problems into satisfiability problems, enabling it to tackle a range of reasoning problems, such as SAT and GCP.
@@
We evaluate the performance of our approach on three constraint satisfaction problems: logical deduction, Boolean satisfiability, and graph coloring. Our analysis yields two key findings: firstly, on simple artificial problem instances, DiLA boosts the inference accuracy of LLMs to 100% and consistently outperforms solver-aided approaches with improved runtime. Secondly, for real-world problems that current solvers struggle with, DiLA showcases robustness and remarkable efficiency in handling these complex test cases, thereby opening up opportunities for further real-world applications.
@@ 2 Motivation
Figure 2: Exemplar comparison of solving graph coloring problems by different approaches. Direct prompts by GPT-4 make errors when generating the color assignment step by step; LOGIC-LM based on the Z3 symbolic solver cannot solve parsed constraints and outputs unknown (both errors are highlighted in red). In contrast, DiLA can generate the correct answer by combining the strengths of LLMs and the differential logic layer.
This paper explores the potential of language understanding and logical reasoning capabilities in LLMs. Traditionally, the prevailing approaches have either relied solely on LLMs for step-by-step reasoning (Wang et al., 2023; Zhou et al., 2022) or offloaded reasoning tasks to off-the-shelf solvers (Olausson et al., 2023; Ye X, 2023). However, we propose a third approach, as these two extremes either underutilize or over-rely on LLMs during reasoning. Specifically, DiLA leverages the powerful understanding abilities of LLMs to extract logical formulas and generate possible solutions. By utilizing LLMs, such as GPT-4, as a solution generation engine, we can produce an initial solution based on input semantic constraints, laying the groundwork for further refinement.
@@
Consequently, our objective is to identify a synergistic approach that combines the strengths of LLMs and a differential logic layer, thereby circumventing the limitations of traditional solvers. This approach leverages the LLM’s capacity to comprehend logical formulas while concurrently utilizing the logic layer’s refinement abilities to achieve accurate solutions. Figure 2 illustrates a comparison between CoT, LOGIC-LM, and our proposed DiLA. The LLM alone may introduce logical flaws during step-by-step inference, such as assigning the same color to vertex 1 and vertex 4 despite their edge connection, and LOGIC-LM may struggle with complex reasoning problems due to its backbone solver’s limitations. In contrast, DiLA produces accurate answers through the collaboration of the LLM and the differential logic layer. Specifically, DiLA first uses the LLM to parse a natural language input into logic constraints and generate an initial solution based on its semantic understanding, then employs the differential logic layer to refine this initial solution.
3 Differenial Logic Layer-Aided Language Models
3.1 Overview
@@
In this section, we present DiLA, which augments LLM with the ability of logical reasoning by incorporating a differential logic layer. More specifically, DiLA addresses the challenge of using LLMs to tackle canonical reasoning tasks expressed in natural language. These tasks typically involve presenting a set of premises and constraints, prompting questions that necessitate intricate deductive reasoning over the provided inputs, which remains a formidable challenge even for contemporary LLMs (Valmeekam et al., 2022).
@@ 3.2 Problem Formulator and Initialization
Intuitively, LLMs may struggle with directly solving complex reasoning problems. However, they have demonstrated a notable ability to comprehend textual inputs and translate them into formal programs, such as mathematical equations (He-Yueya et al., 2023) or satisfiability modulos (Ye X, 2023). Notably, the SAT problem can serve as a versatile intermediate step for solving a broad range of constraint satisfaction problems, provided they can be expressed using Boolean variables. Specifically, problem instances from NP-complete domains, such as Graph Coloring and Set Cover, can be seamlessly encoded into SAT problem specifications, thereby allowing for efficient solutions via SAT algorithms (Stechly et al., 2024; Liu et al., 2023). Therefore, we harness SAT encoding as a general bridge to tackle these constraint satisfaction problems in practical settings.
Specifically, given a problem description in natural language, DiLA prompts an LLM with detailed instructions to generate the SAT specification, which includes a set of premises and constraints. Typically, the SAT specification here involves conjunctive normal forms (CNFs), denoted as $\phi(v_{1},\dots,v_{n})$, which is a conjunction of clauses (constraints) $C$. Formally, rules are written in the conjunctive form of clauses $C_{1}\wedge C_{2}\cdots\wedge C_{m}$, which each $C_{j}$ is a constraint. The total rule is satisfied if and only if all of the clauses $C_{1},C_{2},\dots,C_{m}$ are simultaneously True. Each clause represents a disjunction of literals, where a literal is either a propositional variable $v_{i}$ or its complement $\neg v_{i}$, for example, $v_{1}\vee\neg v_{2}$. In other words, if at least one literal in a clause is True, this clause would also be True. Variables can be assigned logic values, either $1$ or $-1$, representing True or False, respectively  [^1] In other works, they may claim the logic value of each literal is 00 or $1$. It should be noted that the two claims are equal under simple mathematical transformations.. See Appendix A for more details on the SAT specification.
Aside from problem formulation, leveraging an LLM can be a valuable strategy to generate an initial solution. We observe that, after generating the SAT specification, we can prompt LLMs, like GPT-4 (Achiam et al., 2023) or Llama-3 (Meta, 2024), to produce a potential solution or a set of possible solutions. This can be achieved by framing the problem as a natural language query, such as “What is the logical solution based on the premises?” or “Can you provide a possible answer from these given constraints?” (see Appendix B for more details). The LLM’s response can then serve as a starting point for further refinement and validation, allowing us to build upon its output and iteratively improve the solution through logical analysis and reasoning. By harnessing the LLM’s ability to understand semantic constraints, we can tap into its potential to facilitate the initial solution-finding process and accelerate solving progress towards a well-reasoned answer.
@@ 3.3 From SAT to Differential MaxSAT Formulations
In fact, SAT is very efficient at expressing constraint satisfaction problems in that many of the standard NP-complete problems, like graph coloring, can be straightforwardly formulated and efficiently solved as SAT problems (Buss and Nordström, 2021; Liu et al., 2023). Traditionally, a SAT solver, such as Z3 and Kissat, is leveraged to determine a satisfying assignment for the given constraint formula $\phi$. Contemporary SAT solvers are founded on the Conflict-Driven-Clause-Learning (CDCL) algorithm, which excels in its ability to learn from conflicts and use that conflict knowledge to prune the branch-and-bound search space more effectively. However, existing CDCL-based SAT solvers still suffer from exponential searching space and are unable to correct errors through a learning-from- mistakes system, resulting in an infinite loop in solving complex SAT problems (Shi et al., 2023).
In this study, when an NL reasoning problem is fed into DiLA, our goal is to determine its solution that can satisfy all logical constraints. To achieve this, we might incorporate a solver as an additional logic reasoner for the LLM, as in SATLM (Ye X, 2023) or LOGIC-LM (Pan et al., 2023). However, the exponential search complexity inherent in heuristic-based SAT solvers poses a significant challenge, limiting their effectiveness when dealing with complex real-world problems. Therefore, a key issue is how to design an efficient solver surrogate that can both be seamlessly integrated into LLMs and efficiently address logical reasoning problems.
@@
where $\bigwedge$ is the logical “and” symbol. Indeed, the objective value in Equation 2 is 0 if and only if a satisfiable solution can be found. Our goal is to establish a continuous upper bound, referred to as the “loss”, for each clause to quantify its level of unsatisfiability. In essence, the loss takes an upper bound if the clause is unsatisfied, and by minimizing this loss, we can strive to push it closer to satisfaction. To make a purely quadratic loss function as in  (Wang and Kolter, 2019), we introduce $v_{0}=1$ and $s_{0j}=-1$ in Equation 2. Therefore, the minimization problem in Equation 2 can be solved by transforming into a quadratic loss function as:
@@
Figure 4: Examplar illustration of VCG.
@@ 3.4 Differential Logic Layer
We envision the logic layer being used primarily at the top of LLMs, embedding logical formulas $\phi$ produced by the backbone LLM, taking LLM-initialized variable assignments as inputs, and producing outputs that are consistent with $\phi$. Specifically, we draw an analogy between the fully connected (FC) layer and the variable-clause graph (VCG), as depicted in Figure 4, where the weight for positive literals in the clause is 1 (solid arrow) and negative literals in the clause is -1 (dashed arrow). Specifically, we map each variable to an input neuron in an FC layer, each clause to an output neuron, and the coefficients to the weights in the linear transformation with a zero bias vector. While this analogy holds, there are two crucial differences between traditional FC layers and our proposed logic layer. Firstly, the differential logic layer has no unknown parameters, whereas FC layers require data-driven training to learn their weights. Secondly, each clause in the logic layer is only partially connected to variables, in contrast to fully connected layers, where all input neurons are connected to all output neurons.
@@
Algorithm 1 The forward pass of logic layer
0:  Solution $\boldsymbol{v}^{k}\in\mathbb{R}^{n}$ at $k$-th epoch.
0:  $y^{k}$, final solution $\boldsymbol{v}^{*}$.
1:  $\tilde{\boldsymbol{v}}^{k}\leftarrow[\boldsymbol{v}^{k}_{i}>0:i=1,\dots,n]$;
2:  $\phi^{\prime}\leftarrow\phi(\tilde{v}^{k}_{1},\dots,\tilde{v}^{k}_{n})$;
3:  if $\phi^{\prime}$ is satisfiable then
4:     $y^{k}\leftarrow$ True;
5:     $\boldsymbol{v}^{*}\leftarrow\tilde{\boldsymbol{v}}^{k}$ ;
6:  else
7:     $y^{k}\leftarrow$ False;
8:  end if
Backward Pass. The backward pass is responsible for computing the gradients of the layer inputs and derives updates to variables that steer towards satisfying the constraints $\phi$. A crucial aspect of the backward pass is identifying the input variables that contribute most to the unsatisfiability of the constraint formulas. It is well-established that variables in $\bar{I}$ form the unsatisfiable subset and are, therefore, more likely to be sources of conflict. Conversely, variables not present in $\bar{I}$ can have their gradients set to zero, as their absence in the conflict clauses provides no evidence regarding the correctness or incorrectness of their values. Inspired by the stochastic local search (SLS) algorithm, commonly used in constraint satisfaction problems (Chu et al., 2023), we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value at each iteration. However, our logic layer diverges from SLS in that it employs a “differential” variable selection mechanism during backpropagation, whereas SLS relies on meta-heuristics.